Compute the ending inventory for Dino Radja Company for 2017 through 2022 using the dollar-value LIFO method. E8.23 (LO 4) (Dollar-Value LIFO) The following information relates to the Jimmy Johnson Company.    

Date       Ending Inventory (End-of-Year Prices)    Price Index

December 31, 2016 $ 70,000 100

December 31, 2017    90,300 105

December 31, 2018    95,120 116

December 31, 2019   105,600 120

December 31, 2020   100,000 125

Instructions Use the dollar-value LIFO method to compute the ending inventory for Johnson Company for 2016 through 2020.
Answer #1

Calculate ending inventory

Ending inventory
2016 70000
2017 90300/1.05 = 86000 (70000+16000*1.05) = 86800
2018 95120/1.16 = 82000 (70000+12000*1.05 = 82600
2019 105600/1.20 = 88000 (82600+6000*1.20) = 89800
2020 100000/1.25 = 80000 (70000+10000*1.05) = 80500